OpenStudy (anonymous):

Solve using substitution:5y − x = -2 -2x = 6 − 10y No solution Infinitely many solutions (3, -1) (-8, -2)

OpenStudy (anonymous):

@surjithayer help please

OpenStudy (anonymous):

@AlwaystheBookworm

OpenStudy (anonymous):

5y-x=-2 -2x=6-10y Change each equation to x+y=b -x+5y=-2 -2x+10y=6 Multiply the top equation by -2. 2x-10y=4 -2x+10y=6 Cancel out the x. -10y=4 10y=6 Add ys and constants. 0=10 No solutions

OpenStudy (anonymous):

thank you